(Only question 21, but you need the information in 20 to answer)

20.    ll How much work is required to completely separate two charges (each -1.4 micro C) and leave them at rest if they were initially 8.0 mm apart? 2.2 J

21.        ll In Exercise 20, if the two charges are released at their initial separation distance, how much kinetic energy would each have when they are very distant from one another?
Answer #1

As in question 20 the work is required to completely separate two charges (each -1.4 micro C) and leave them at rest if they were initially 8.0 mm apart is 2.2J

so,

At their initial separation distance the kinetic energy = work /2 = 2.2/2 =1.1 j

Answer : 1.1J